# PFML: Self-Supervised Learning of Time-Series Data Without Representation Collapse

### v1

@@ 4.5 Additional Hyperparameter Experiments
For each data modality, we also experimented with different configurations of masking probability $p_{\text{m}}$ and the length of the masks $l_{\text{m}}$. We ran PFML pre-training using different configurations of $p_{\text{m}}$ and $l_{\text{m}}$, and then we fine-tuned the pre-trained models. For IMU and speech data, we only experimented with one classification task each, namely classification of movement from IMU data and classification of valence from speech data. The results for different configurations of $p_{\text{m}}$ and $l_{\text{m}}$ for IMU, speech, and EEG data are shown in Appendix C in Tables 7, 8, and 9, respectively. For IMU data, the differences between different masking strategies are rather small, whereas for speech and EEG data the selection of masking hyperparameters has a notable effect on fine-tuning performance.
@@ 5 Conclusion
In this paper, we presented PFML, a novel SSL algorithm for time-series data that avoids the common SSL issue of representation collapse. PFML operates by predicting statistical functionals of the input signal corresponding to masked embeddings, given a sequence of unmasked embeddings. We demonstrated the effectiveness of PFML using five different classification tasks across three different data modalities: infant posture and movement classification from multi-sensor IMU data, emotion recognition from speech data, and sleep stage classification from EEG data. Our results show that PFML is superior to a conceptually similar SSL method, MAE. Our results also show that PFML is competitive against the current state-of-the-art data modality agnostic SSL method, data2vec, while being conceptually simpler and without suffering from representation collapse. The fact that PFML matches the performance of data2vec while also avoiding the issue of representation collapse renders PFML more straightforward to apply to new time-series data domains, such as in the case of clinical time-series data. The present work may also be extended to other domains than time-series data, such as images where functionals could be computed of, e.g., image patches.
@@
We selected the present set of 11 functionals for their effectiveness across the three data modalities used in the present study, aiming for potential generalizability and a robust starting point to other data domains and downstream tasks. However, carefully selecting the number and type of functionals specifically for different modalities may lead to better results than presented here. Also, we did not include data augmentation in our pre-training processes to save computational time for PFML pre-training, as we wanted to pre-compute the functionals before the model training. As shown in e.g. Chen et al. (2020); Grill et al. (2020); He et al. (2022); Balestriero et al. (2023), data augmentation during pre-training may lead to improved performance on downstream tasks. Nonetheless, performing masking for randomly sampled frames is already a form of data augmentation in itself. Furthermore, other model architectures besides CNN-based encoders or Transformer encoder blocks could also be used, and this may improve PFML pre-training performance. Lastly, we acknowledge that typically SSL pre-training is run with very large minibatch sizes using multiple GPUs, and the results of the present experiments might improve with larger minibatch sizes. However, to promote reproducibility and encourage other researchers to try PFML, we deliberately pre-trained our models using relatively small minibatches so that the pre-training processes could be run on a single GPU with 16 GB of VRAM. As detailed in Appendix D, our method used only a moderate amount of computational resources.
@@ Appendix C Results on Additional Hyperparameter Experiments
This section provides the result tables for the additional hyperparameter experiments in Section 4.5. Table 6 presents the fine-tuning results for the comparison between masking either the model inputs or embeddings during PFML pre-training. Tables 7, 8, and 9 show the fine-tuning results for different configurations of masking probabilities ($p_{\text{m}}$) and mask lengths ($l_{\text{m}}$) for multi-sensor IMU, speech, and EEG data, respectively. Table 10 shows the fine-tuning results for discarding some of the 11 functionals in PFML pre-training for multi-sensor IMU data. Finally, Table 11 presents the fine-tuning results for different mask types for PFML pre-training for multi-sensor IMU data.
@@
Appendix D Additional Information on Computational Resources
@@
The research project required more computations than what is reported in the present paper: For each data modality and pre-training algorithm, preliminary experiments were conducted in order to find suitable hyperparameters for both the pre-training and fine-tuning processes (Tables 4 and 5, respectively).

### v2

@@ 6 Additional Hyperparameter Experiments
For each data modality, we also experimented with different configurations of masking probability $p_{\text{m}}$ and the length of the masks $l_{\text{m}}$. We ran PFML pre-training using different configurations of $p_{\text{m}}$ and $l_{\text{m}}$, and then we fine-tuned the pre-trained models. For IMU and speech data, we only experimented with one classification task each, namely classification of movement from IMU data and classification of valence from speech data. The results for different configurations of $p_{\text{m}}$ and $l_{\text{m}}$ for IMU, speech, and EEG data are shown in Appendix C in Tables 8, 9, and 10, respectively. For IMU data, the differences between different masking strategies are rather small, whereas for speech and EEG data the selection of masking hyperparameters has a notable effect on fine-tuning performance.
@@ 7 Conclusion
In this paper, we presented PFML, a novel SSL algorithm for time-series data that avoids the common SSL issue of representation collapse. PFML operates by predicting statistical functionals of the input signal corresponding to masked embeddings, given a sequence of unmasked embeddings. We demonstrated the effectiveness of PFML using five different classification tasks across three different data modalities: infant posture and movement classification from multi-sensor IMU data, emotion recognition from speech data, and sleep stage classification from EEG data. Our results show that PFML is superior to both a conceptually similar SSL method, MAE, and a contrastive learning-based SSL method, TS2Vec. Our results also show that PFML is on par with the current state-of-the-art data modality agnostic SSL method, data2vec, while being conceptually simpler and without suffering from representation collapse. The pros and cons for each of the SSL methods used in the present study are shown in Table 4. The fact that PFML matches the performance of data2vec while also avoiding the issue of representation collapse renders PFML more straightforward to apply to new time-series data domains, such as in the case of clinical time-series data. The present work may also be extended to other domains than time-series data, such as images where functionals could be computed of, e.g., image patches.
Table 4: A listing of pros and cons for each of the SSL methods used in the present experiments. For further details on these results, refer to Appendix D.
@@
We selected the present set of 11 functionals for their effectiveness across the three data modalities used in the present study, aiming for potential generalizability and a robust starting point to other data domains and downstream tasks. However, carefully selecting the number and type of functionals specifically for different modalities may lead to better results than presented here. Also, we did not include data augmentation in our pre-training processes to save computational time for PFML pre-training, as we wanted to pre-compute the functionals before the model training. As shown in e.g. Chen et al. (2020); Grill et al. (2020); He et al. (2022); Balestriero et al. (2023), data augmentation during pre-training may lead to improved performance on downstream tasks. Nonetheless, performing masking for randomly sampled frames is already a form of data augmentation in itself. Furthermore, other model architectures besides CNN-based encoders or Transformer encoder blocks could also be used, and this may improve PFML pre-training performance. Lastly, we acknowledge that typically SSL pre-training is run with very large minibatch sizes using multiple GPUs, and the results of the present experiments might improve with larger minibatch sizes. However, to promote reproducibility and encourage other researchers to try PFML, we deliberately pre-trained our models using relatively small minibatches so that the pre-training processes could be run on a single GPU with 16 GB of VRAM. As detailed in Appendix E, our method used only a moderate amount of computational resources.
@@ Appendix C Results on Additional Hyperparameter Experiments
This section provides the result tables for the additional hyperparameter experiments in Section 6. Table 7 presents the fine-tuning results for the comparison between masking either the model inputs or embeddings during PFML pre-training. Tables 8, 9, and 10 show the fine-tuning results for different configurations of masking probabilities ($p_{\text{m}}$) and mask lengths ($l_{\text{m}}$) for multi-sensor IMU, speech, and EEG data, respectively. Table 11 shows the fine-tuning results for discarding some of the 11 functionals in PFML pre-training for multi-sensor IMU data. Finally, Table 12 presents the fine-tuning results for different mask types for PFML pre-training for multi-sensor IMU data.
@@
Appendix D Further Details on the Pros and Cons of Each SSL Method
This section provides further details on the results presented in Table 4. Table 13 presents the numerical values behind the results of Table 4.
Table 13: A listing of pros and cons for each of the SSL methods used in the present experiments. Up arrow ($\uparrow$) indicates that a higher score is better, whereas a down arrow ($\downarrow$) indicates that a lower score is better.
The values for the columns in Table 13 have been defined as follows:
  • End-use performance: These results are the average of the UAF1 (%) and UAR (%) values in Table 1 across all five classification tasks.
  • Robust to representation collapse: This is a “Yes/No” decision based on Table 3.
  • Expected pre-training time: The expected pre-training time $t_{\text{pt}}(P_{\text{m}})$ for pre-training method $P_{\text{m}}$ is computed as
    $t_{\text{pt}}(P_{\text{m}})=\left(avg\_num\_collapses+1\right)\cdot avg\_training\_time\ ,$ (5)
    where $avg\_num\_collapses$ is the average number of representation collapses across all three data modalities (Table 3) and $avg\_training\_time$ is the average pre-training time (in hours) across all three data modalities. Note that for MAE, we considered the average number of representation collapses to be zero, since there was only a single representation collapse which was most probably due to bad luck in model weight initialization. Also note that model training time can vary significantly depending on the amount and type of training data, the software implementation, and the computational hardware used. However, when taking the expected number of representation collapses during pre-training into account, the SSL methods used in the present experiments can be categorized into three distinct groups.
  • Flexibility of neural network architecture: This is a “Yes/No” decision based on whether the SSL algorithm has a free choice of the encoder and classifier architecture. In TS2Vec, the pre-training algorithm is built into the pre-defined CNN encoder architecture, and therefore it is not possible select the architecture of the encoder freely.
  • Linear separability of SSL features: These results are the average of the UAF1 (%) and UAR (%) values in Table 2 across all five classification tasks.
  • Number of hyperparameters: These values are a simple count of the number of different SSL algorithm-specific hyperparameters that affect the performance of the SSL algorithm. These algorithm-specific hyperparameters for each SSL algorithm are:
    • TS2Vec: 1) Minibatch size, 2) Number of negative samples in minibatch, 3) Data augmentation method
    • MAE: 1) Masking probability, 2) Mask length
    • data2vec: 1) Masking probability, 2) Mask length, 3) Student network learning rate, 4) Teacher network, initial weight update rate, 5) Teacher network, final weight update rate, 6) Teacher network, number of weight update rate transitions, 7) Training targets
    • PFML: 1) Masking probability, 2) Mask length, 3) Functionals
Appendix E Additional Information on Computational Resources
@@
The research project required more computations than what is reported in the present paper: For each data modality and pre-training algorithm, preliminary experiments were conducted in order to find suitable hyperparameters for both the pre-training and fine-tuning processes (Tables 5 and 6, respectively).
Appendix F Additional TS2Vec Results
This section provides all TS2Vec-related results of the present experiments. Table 15 shows a summary of all of the results regarding TS2Vec for IMU, speech, and EEG data, further described in Section 4.5.
Table 15: All TS2Vec-related results across the three different data modalities (IMU, speech, and EEG data). Starting from the top, the table includes results for model fine-tuning with a randomly initialized model consisting of the TS2Vec encoder and Transformer classifier, followed by results for the same model with the encoder pre-trained using TS2Vec. The rest of the results are without the Transformer: First, linear evaluation results for the TS2Vec encoder without pre-training, followed by linear evaluation results using a TS2Vec pre-trained encoder. Finally, following the procedure used in the original TS2Vec paper (Yue et al., 2022), results are provided using an SVM classifier with an RBF kernel for the SSL features obtained from the pre-trained TS2Vec encoder.